# Emergence of the Primacy Effect in Structured State-Space Models

### v4

@@ Abstract
Human and animal memory for sequentially presented items is well-documented to be more accurate for those at the beginning and end of the sequence, phenomena known as the primacy and recency effects, respectively. By contrast, artificial neural network (ANN) models are typically designed with a memory that decays monotonically over time. Accordingly, ANNs are expected to show the recency effect but not the primacy effect. Contrary to this theoretical expectation, however, the present study reveals a counterintuitive finding: a recently developed ANN architecture, called structured state-space models, exhibits the primacy effect when trained and evaluated on a synthetic task that mirrors psychological memory experiments. Given that this model was originally designed for recovering neuronal activity patterns observed in biological brains, this result provides a novel perspective on the psychological primacy effect while also posing a non-trivial puzzle for the current theories in machine learning.
@@ 1 Introduction
Human and animal memory for sequentially presented items is well-documented to be more accurate for those appearing at the beginning and end of the sequence—phenomena known as the primacy and recency effects, respectively (Ebbinghaus,, 1913; Murdock,, 1962; Glanzer and Cunitz,, 1966). For example, when a sequence of random integers such as $49,75,\dots,5,38$ is presented in that order, the initial ($49,75$) and final ($5,38$) items are more likely to be recalled accurately at the end of the presentation.
By contrast, artificial neural network (ANN) models are typically designed with a memory that decays monotonically over time (Bengio et al.,, 1994; Jaeger,, 2001; Jaeger and Haas,, 2004). Thus, ANNs are expected to show the recency effect but not the primacy effect.
Contrary to this theoretical expectation, however, the present study reveals a counterintuitive finding: a recently developed ANN architecture—called structured state-space models (Gu et al.,, 2020; Gu et al., 2022b, )—exhibits the primacy effect when trained and evaluated on a synthetic task that mirrors psychological memory experiments. Given that this model was originally designed for recovering neuronal activity patterns observed in biological brains, this result provides a novel perspective on the psychological primacy effect while also posing a non-trivial puzzle for the current theories in machine learning.
The remainder of this paper is organized as follows. The next section first reviews ANN models for time-series processing. After this preliminary discussion, the Methods section details the task and model specifications of the present study, and the subsequent section presents the results. Finally, the Discussions section summarizes the findings and discusses their implications in relation to previous research.
2 Preliminaries on ANNs for Time-Series Processing
2.1 Recurrent Neural Networks
The traditional ANN-based approach to time-series processing relies on recurrent neural networks (RNNs; Elman,, 1990). RNNs sequentially process inputs, updating their internal state (represented by a real-valued vector) at each time step based on both the current input and the previous state. The output is then generated through a feedforward transformation of this internal state at each step. In other words, RNNs map input time series to target output sequences through the latent dynamics of their internal state.
Theoretically, RNNs possess Turing-complete computational power, meaning they can simulate an arbitrary computational model in an “idealized” setting (Siegelmann and Sontag,, 1992, 1995; Siegelmann,, 1999). This theoretical capability has driven widespread applications of RNNs in domains such as language modeling (Sundermeyer et al.,, 2012; Graves,, 2013), speech synthesis (Kalchbrenner et al.,, 2018), and weather nowcasting (Shi et al.,, 2015).
However, the “idealized” computational properties of RNNs are unattainable in empirical implementations. For instance, when deployed on standard computers, the weight parameters of RNNs can only be represented with finite precision (Chen et al.,, 2018; Weiss et al.,, 2018), preventing them from storing an unbounded number of observations. Additionally, RNNs exhibit memory decay over time, which has led researchers to explore methods for extending their memory retention capabilities (Hochreiter and Schmidhuber,, 1997; Arjovsky et al.,, 2016; Neil et al.,, 2016; Chang et al.,, 2017; Jing et al.,, 2017, 2019).
@@ 2.2 Structured State-Space Models
Recently, researchers have found that continuous-time models can achieve more persistent memory retention than discrete-time RNNs (Zhang et al.,, 2018; Voelker et al.,, 2019; Gu et al.,, 2020). The fundamental principle underlying these models is the polynomial approximation of observed signals (called the HiPPO framework; Gu et al.,, 2020). Specifically, given a single-channel input signal $x(\cdot)$, represented as a function of time, its approximation up to a given time point $t$ can be approximated by a linear combination of polynomials:
$x|_{\leq t}(\cdot)\approx\sum_{n=0}^{N-1}h_{n}(t)P_{n}(\cdot)$
where $P_{n}$ denotes the basis polynomial of degree $n$ and $h_{n}(t)$ are the optimal coefficients for the approximation at time $t$111 The optimal coefficients can be determined as $h_{n}(t)=\langle x|_{\leq t},P_{n}\rangle=\int x|_{\leq t}(s)P_{n}(s)d\mu^{(t)}(s)$ when $\{P_{n}\}_{n=0}^{N-1}$ form an orthogonal basis with respect to the time-dependent measure $d\mu^{(t)}(\cdot)$. Then, these coefficients offer a finite- and constant-dimensional representation of the input signal up to time $t$. The framework can be naturally extended to multi-channel signals by performing channel-wise approximations in parallel, yielding $h_{n}^{(m)}(t)$ for each channel $m$.
@@
where $\mathbf{h}(t):=(h_{0}(t),\dots,h_{N-1}(t))^{\mathsf{T}}$, and $A$ and $B$ are matrices of size $N\times N$ and $N\times 1$, respectively. The values on $A$ and $B$—referred to as the state and input matrices, respectively—depend on the choice of the underlying polynomial basis, and can also be adjusted via gradient-based optimization. A feedforward transformation of the state vector $\mathbf{h}(t)$ (achieved via left-multiplication by another matrix $C$) yields a (possibly multi-channel) output signal $\mathbf{y}(t)=C\mathbf{h}(t)\in\mathbb{R}^{M}$.333 The general formulation of the SSM incorporates an additional matrix $D$, which establishes a direct feedforward connection between the input and output signals, expressed as $\mathbf{y}(t)=C\mathbf{h}(t)+D\mathbf{x}(t)$. In practice, however, $D$ is often set to the identity matrix, effectively reducing the feedforward transformation to a simple residual connection (He et al.,, 2016). The resulting mapping $x\mapsto\mathbf{y}$ is termed the structured state-space model (hereinafter abbreviated as SSM; Gu et al.,, 2021; Gu et al., 2022b, ).
@@
$\bar{A}:=$ $\left(I-\frac{\Delta t}{2}A\right)^{-1}\left(I+\frac{\Delta t}{2}A\right)$ $\bar{B}:=$ $\left(I-\frac{\Delta t}{2}A\right)^{-1}\Delta tB$
@@
where $\Delta t:=t_{j+1}-t_{j}$ ($\forall j=1,\dots,L-1$) defines the time-step size. This time-step parameter is treated as learnable, allowing the model to automatically adjust the time scale of its state-space dynamics to align with that of the input signal. Moreover, in multi-channel settings, the model can represent multi-scale dynamics by assigning distinct $\Delta t^{(m)}$ values to each channel $m$.
@@
This flexibility extends the applicability of SSMs to inherently discrete data lacking overt continuous dynamics (e.g., text languages; Gu and Dao,, 2024; Dao and Gu,, 2024); the model can jointly lean latent representations (or embeddings) of discrete inputs along with their (pseudo-)continuous dynamics via gradient-based optimization. Furthermore, SSMs can be hierarchically stacked to construct deeper and more expressive ANNs, where each layer processes latent signals received from lower layers.
@@
Figure 1: Illustration of the continuous (left; $K(\tau):=e^{\tau A}B$) and discretized (right; $\bar{K}_{j}:=\bar{A}^{j}\bar{B}$) SSM kernels based on Legendre polynomials with an exponentially decaying measure (Gu et al.,, 2023). To aid intuitive understanding (particularly for readers unfamiliar with convolutional operations), the horizontal axis has been flipped, so that kernel values multiplied with past inputs appear on the left (unlike in the standard visualization of convolutional kernels, where they are placed on the right). The distinct line colors represent selected entries of the kernel vectors, $K_{N}(\tau)$ and $\bar{K}_{j,n}$, each corresponding the approximating polynomial of degree $n\in\{2,3,8,15,32,63\}$. The kernels were evaluated at $\tau=j\Delta t$ for $j=0,\dots,255$ and $\Delta t\in\{0.001,0.01,0.1,1.0\}$. Increasing $\Delta t$ results in growing discrepancies between the continuous and discretized kernels.
Previous studies have identified that the time-step size, $\Delta t$, as a critical factor in determining the success/failure of SSMs. Intuitively, a small $\Delta t$ results in minor state updates, yielding slow dynamics in the state space, $\mathbf{h}(t_{j+1})-\mathbf{h}(t_{j})$; conversely, a large $\Delta t$ induces rapid state transitions (Gu et al.,, 2021; Gu and Dao,, 2024). As a consequence, $\Delta t$ governs the memory decay properties of the SSM; although the models assume an exponentially decaying measure in continuous time (particularly when employing Legendre/Laguerre polynomials; Gu et al.,, 2023), choosing small $\Delta t$ values can allocate relatively large weights to input samples from distant time steps, thereby compromising single-step discriminability, which is better preserved with larger $\Delta t$ (see left panels in Fig. 1).
It is also important to note that discretization methods are generally designed under the assumption of sufficiently small $\Delta t$, for effectively approximating the limit $\Delta t\to 0$. Consequently, setting $\Delta t$ too large introduces discrepancies between the continuous and discretized dynamics (as illustrated by the contrast between the left vs. right panels in Fig. 1).
2.3 Finite-Queue/Markov Models
Finally, it is worth mentioning finite-queue/Markov models, which encompass the current goldstandard architecture of ANNs, Transformer (Vaswani et al.,, 2017). These models are equipped with an explicit storage mechanism that retains raw observation records up to a prespecified limit. Formally, they enable arbitrary computations at a given discrete time $j$ while maintaining full access to the most recent $M$ observations, $x_{j-M+1},\dots,x_{j}$. At each time step, the oldest observation ($x_{j-M}$) is discarded as new data enter the storage buffer.
By design, finite-queue/Markov models do not exhibit the primacy effect once earlier observations inputs fall outside the accessible memory window. Instead, they align with the classical view of biological short-term memory Miller, (1956), which has traditionally been linked with the recency effect (Waugh and Norman,, 1965; Glanzer and Cunitz,, 1966; Atkinson and Shiffrin,, 1968).
Nonetheless, given sufficient computational resources, the memory capacity of modern models like Transformers can cover most empirical sequences, rendering the out-of-window effects negligible in practice. In such cases, models with naive implementations assign equal importance to all observations. Most notably, the canonical Transformer is inherently agnostic to input order in the absence of positional encoding (Vaswani et al.,, 2017), implying a form of bias-free memorization. Nevertheless, as discussed in the following section, even such unbiased models can exhibit the primacy effect when trained on specific datasets or tasks.
@@ 2.4 Data-Driven Primacy Effect in Language Models
Several prior studies have documented the primacy effect of ANNs trained on the autoregressive language modeling task. Wang et al., (2023) investigated positional biases in a Transformer-based large language model (LLM) using a prompting-based approach. Specifically, a list of action or event labels was sequentially presented (e.g., “Label 1: change_pin”, “Label 2: card_arrival”, “Label 3: activate_my_card”). The model was then given a query prompt specifying a target action/event (e.g., “Target Text: I need a new PIN.”) along with an instruction statement (e.g., “Which label matches the intent expressed in the Target Text?”). The primacy effect was observed as a greater frequency of the initially presented labels in the model’s responses.444 According to the published code, the indices assigned to the labels (e.g., “2” in “Label 2: card_arrival”) consistently corresponded to their order of presentation; that is, “Label $j$: xxx” was never listed before “Label $j^{\prime}$: yyy” if $j^{\prime}<j$, even though the label contents (“xxx” and “yyy”) were disassociated from the label indices and ordered randomly. Hence, the reported primacy effect could be attributed to a bias in model responses toward lower-indexed labels, which appear more frequently than larger integers (e.g., 2 vs. 67), rather than to greater attentions to initially presented labels. Comparable findings have been reported across different LLM implementations and benchmark datasets (Eicher and Irgolič,, 2024; Guo and Vosoughi,, 2024; Janik,, 2024; Liu et al.,, 2024).
@@
In contrast to prior investigations, the present study examines the emergence of the primacy effect in ANNs preventing the inheritance of human-induced bias. Specifically, the models are trained on a synthetic memorization task designed based on psychological experiments conducted with humans and other animals (Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985). The following section details the task formulation and the model architecture.
3 Methods
3.1 Task
The memorization patterns of ANNs were assessed using the binary memory verification task (Fig. 2; a.k.a. serial probe recognition; Wickelgren and Norman,, 1966; Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985).555 The most widely adopted task for assessing the primacy effect in human memory is free recall, in which participants are presented with a sequence of items and subsequently asked to recall them in an order-agnostic manner (Murdock,, 1962; Glanzer and Cunitz,, 1966). While this paradigm can be technically formulated as a loss function for ANNs (Cuturi,, 2013), initial explorations of this study revealed that model performance remained suboptimal under this approach, yielding lower accuracy than in theoretically more demanding tasks requiring order-sensitive reconstruction. Consequently, the present study adopted the more machine learning-friendly task based on binary verification. Remarkably, this task has also been used to assess the memory capacity of non-human animals, which are unable to perform free recall (Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985). In this task, the models were first presented with a sequence of randomly generated, non-repeating integers (hereinafter referred to as study items). Subsequently, they received another sequence of integer queries and were trained to determine whether each query token was present (labeled as 1) or absent (labeled as 0) in the study items. To construct these queries, the study items were first shuffled, and then, with a probability of $p=0.5$, each shuffled token was replaced with a randomly sampled integer from the complement set of the study items (termed distractors).
@@ 3.1 Task
To build test sequences, the held-out study items were randomly ordered, and queries were generated by first shuffling and then cyclically shifting them (e.g., $(2,8,11,29)\mapsto\{(2,8,11,29),(8,11,29,2),(11,29,2,8),(29,2,8,11)\}$). This design ensured that each study item was queried in all $L$ possible positions. Finally, either the even- or odd-indexed query positions were replaced with random distractors, resulting in a total of $1024\times L\times 2$ test sequences per trial.

### v5

@@ Abstract
Structured state-space models (SSMs) have been developed to offer more persistent memory retention than traditional recurrent neural networks, while maintaining real-time inference capabilities and addressing the time-complexity limitations of Transformers. Despite this intended persistence, the memory mechanism of canonical SSMs is theoretically designed to decay monotonically over time, meaning that more recent inputs are expected to be retained more accurately than earlier ones. Contrary to this theoretical expectation, however, the present study reveals a counterintuitive finding: when trained and evaluated on a synthetic, statistically balanced memorization task, SSMs predominantly preserve the initially presented data in memory. This pattern of memory bias, known as the primacy effect in psychology, presents a non-trivial challenge to the current theoretical understanding of SSMs and opens new avenues for future research.
Figure 1: Graphical abstract of the key finding. When trained on a memorization task with sequentially presented items, the SSM achieves the highest accuracy for the earliest items (orange curve). This contrasts with the theoretical design of the SSM’s memory mechanism, which assigns exponentially diminishing importance to older observations (blue curve).
@@ 1 Introduction
In recent years, structured state-space models (SSMs) have garnered increasing attention as a backbone of next-generation artificial intelligent systems (Gu et al., 2022b, ; Gu and Dao,, 2024). SSMs were developed to provide more persistent memory retention than traditional recurrent neural networks (RNNs), while maintaining real-time inference capabilities and addressing the time-complexity limitations of Transformers.
Despite this intended persistence, the memory mechanism of canonical SSMs is theoretically designed to decay monotonically over time (Gu et al.,, 2023); that is, more recent inputs are expected to be retained more accurately than earlier ones. For example, when a sequence of random integers such as $49,75,\dots,5,38$ is presented in that order, the final items ($5,38$) are theoretically more likely to be recalled accurately at the end of the sequence (blue curve in Figure 1).
Contrary to this theoretical expectation, however, the present study reveals a counterintuitive finding: when trained and evaluated on a synthetic, statistically balanced memorization task, SSMs predominantly preserve the initially presented data (e.g., $49,75$) in memory (orange curve in Figure 1). This memory bias is known as the primacy effect in psychology; human and animal memory for sequentially presented items tends to be more accurate for those appearing at the beginning of the sequence (Ebbinghaus,, 1913; Murdock,, 1962; Glanzer and Cunitz,, 1966).111 Human and animal memory is also known to be more accurate for the most recently observed items—a phenomenon termed the recency effect. The theoretical design of SSMs aligns with this opposite pattern. This finding presents a non-trivial puzzle for existing theories and opens new avenues for research on SSMs.
The remainder of this paper is organized as follows. §2 first reviews the formalization and memory characteristics of SSMs. The section also discusses data-driven primacy effects observed in large language models—which inherit human biases embedded in linguistic data—and contrasts these findings with the statistically balanced setting examined in this study. §3 then details the task and model specifications, followed by results in §4. Finally, §5 summarizes the findings and discusses their implications in the context of prior work.
2 Preliminaries and Related Studies
@@ 2.1 Structured State-Space Models
2.1.1 Formalization
To achieve more persistent memory retention than traditional RNNs, the SSM leverages continuous-time dynamics, replacing the discrete-time framework of RNNs (Zhang et al.,, 2018; Voelker et al.,, 2019; Gu et al.,, 2020). The fundamental principle underlying these models is the polynomial approximation of observed signals (called the HiPPO framework; Gu et al.,, 2020). Specifically, given a single-channel input signal $x(\cdot)$, represented as a function of time, its approximation up to a given time point $t$ can be approximated by a linear combination of polynomials:
$x|_{\leq t}(\cdot)\approx\sum_{n=0}^{N-1}h_{n}(t)P_{n}(\cdot)$ (1)
where $P_{n}$ denotes the basis polynomial of degree $n$ and $\{h_{n}(t)\}_{n=0}^{N-1}$ are the optimal coefficients for the approximation at time $t$. When these polynomials form an orthogonal basis with respect to a time-dependent measure $d\mu^{(t)}(\cdot)$, the optimal coefficients can be determined as:
$h_{n}(t)=\langle x|_{\leq t},P_{n}\rangle=\int x|_{\leq t}(s)P_{n}(s)\,d\mu^{(t)}(s)$ (2)
Then, these coefficients offer a finite- and constant-dimensional representation of the input signal up to time $t$. The framework can be naturally extended to multi-channel signals by performing channel-wise approximations in parallel, yielding $h_{n}^{(m)}(t)$ for each channel $m$.
@@ 2.1.1 Formalization
where $\mathbf{h}(t):=(h_{0}(t),\dots,h_{N-1}(t))^{\mathsf{T}}$, and $A$ and BB—referred to as the state and input matrices, respectively—are of size $N\times N$ and $N\times 1$. The values on $A$ and $B$ depend on the choice of the underlying polynomial basis, and can also be adjusted via gradient-based optimization. A feedforward transformation of the state vector $\mathbf{h}(t)$ (achieved via left-multiplication by another matrix $C$) yields a (possibly multi-channel) output signal $\mathbf{y}(t)=C\mathbf{h}(t)\in\mathbb{R}^{M}$.333 The general formulation of the SSM incorporates an additional matrix $D$, which establishes a direct feedforward connection between the input and output signals, expressed as $\mathbf{y}(t)=C\mathbf{h}(t)+D\mathbf{x}(t)$. In practice, however, $D$ is often set to the identity matrix, effectively reducing the feedforward transformation to a simple residual connection (He et al.,, 2016). The resulting mapping $x\mapsto\mathbf{y}$ defines the SSM (Gu et al.,, 2021; Gu et al., 2022b, ).
@@
$\bar{A}:=$ $\left(I-\frac{\Delta t}{2}A\right)^{-1}\left(I+\frac{\Delta t}{2}A\right)$ $\bar{B}:=$ $\left(I-\frac{\Delta t}{2}A\right)^{-1}\Delta tB$ (5)
@@
where $\Delta t:=t_{j+1}-t_{j}$ ($\forall j=1,\dots,L-1$) defines the time-step size. This time-step parameter is treated as learnable, allowing the model to automatically adjust the time scale of its state-space dynamics to align with that of the input signal.444 Despite the data-driven adjustability of $\Delta t$, it is important to recognize that discretization methods are generally designed under the assumption of sufficiently small values on this parameter, for effectively approximating the limit $\Delta t\to 0$. Consequently, setting $\Delta t$ too large introduces discrepancies between the continuous and discretized dynamics (as illustrated by the contrast between the left vs. right panels in Figure 2). Moreover, in multi-channel settings, the model can represent multi-scale dynamics by assigning distinct $\Delta t^{(m)}$ values to each channel $m$.
@@
This flexibility extends the applicability of SSMs to inherently discrete data lacking overt continuous dynamics (e.g., text languages; Gu and Dao,, 2024; Dao and Gu,, 2024); the model can jointly lean latent representations (or embeddings) of discrete inputs along with their (pseudo-)continuous dynamics via gradient-based optimization. Furthermore, SSMs can be hierarchically stacked to construct deeper and more expressive models, where each layer processes latent signals received from lower layers.
2.1.2 Prior Studies on the Memory Dynamics of SSMs
@@
Figure 2: Illustration of the continuous (left; $K(\tau):=e^{\tau A}B$) and discretized (right; $\bar{K}_{j}:=\bar{A}^{j}\bar{B}$) SSM kernels based on Legendre polynomials with an exponentially decaying measure (Gu et al.,, 2023). To aid intuitive understanding, the horizontal axis has been flipped, so that kernel values multiplied with past inputs appear on the left (unlike in the standard visualization of convolutional kernels, where they are placed on the right). The distinct line colors represent selected entries of the kernel vectors, $K_{N}(\tau)$ and $\bar{K}_{j,n}$, each corresponding to the approximating polynomial of degree $n\in\{2,3,8,15,32,63\}$. The kernels were evaluated at $\tau=j\Delta t$ for $j=0,\dots,255$ and $\Delta t\in\{0.001,0.01,0.1,1.0\}$. Increasing $\Delta t$ results in growing discrepancies between the continuous and discretized kernels.
[…]
Owing to the rich theoretical foundations, previous experimental investigations into the memory capacity of SSMs have remained relatively cursory. In particular, most prior works have only reported time-averaged benchmark scores, without offering a detailed analysis of the temporal patterns in memorized vs. overlooked information (Gu et al.,, 2020, 2021; Gu et al., 2022b, ; Gu et al., 2022a, ; Gupta et al.,, 2022). The present study addresses this underexplored question and reveals a primacy effect in SSMs, which stands in stark contrast to their theoretically prescribed memory decay.
@@ 2.2 Data-Driven Primacy Effect in Language Models
Several prior studies have documented the primacy effect of artificial neural networks trained on the autoregressive language modeling task. Wang et al., (2023) investigated positional biases in a Transformer-based large language model (LLM) using a prompting-based approach. Specifically, a list of action or event labels was sequentially presented (e.g., “Label 1: change_pin”, “Label 2: card_arrival”, “Label 3: activate_my_card”). The model was then given a query prompt specifying a target action/event (e.g., “Target Text: I need a new PIN.”) along with an instruction statement (e.g., “Which label matches the intent expressed in the Target Text?”). The primacy effect was observed as a greater frequency of the initially presented labels in the model’s responses. Comparable findings have been reported across different LLM implementations and benchmark datasets (Eicher and Irgolič,, 2024; Guo and Vosoughi,, 2024; Janik,, 2024; Liu et al.,, 2024).
@@
In contrast to these prior investigations, the present study examines the emergence of the primacy effect in the SSM while preventing the inheritance of human-induced bias. Specifically, the models are trained on a synthetic memorization task designed based on psychological experiments conducted with humans and other animals (Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985). The following section details the task formulation and the model architecture.
3 Methods
3.1 Task
@@
The memorization patterns of the SSM were assessed using the binary memory verification task (Figure 3; a.k.a. serial probe recognition in psychology and ethology; Wickelgren and Norman,, 1966; Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985).555 The most widely adopted task for assessing the primacy effect in human memory is free recall, in which participants are presented with a sequence of items and subsequently asked to recall them in an order-agnostic manner (Murdock,, 1962; Glanzer and Cunitz,, 1966). While this paradigm can be technically formulated as a loss function—under the framework of the optimal transport (Cuturi,, 2013)—initial explorations of this study revealed that model performance remained suboptimal under this approach, yielding lower accuracy than in theoretically more demanding tasks requiring order-sensitive reconstruction. Consequently, the present study adopted the more machine learning-friendly task based on binary verification. Remarkably, this task has also been used to assess the memory capacity of non-human animals, which are unable to perform free recall (Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985). In this task, the models were first presented with a sequence of randomly generated, non-repeating integers (hereinafter referred to as study items). Subsequently, they received another sequence of integer queries and were trained to determine whether each query token was present (labeled as 1) or absent (labeled as 0) in the study items. To construct these queries, the study items were first shuffled, and then, with a probability of $p=0.5$, each shuffled token was replaced with a randomly sampled integer from the complement set of the study items (termed distractors).666 An anonymous reviewer noted that the adopted memorization task introduces a non-uniform distribution of relative distances between study items and queries. Specifically, the model has a lower probability of encountering distances of length $L\pm\alpha$, where $L$ is the number of study items, as $\alpha$ increases from 0 to L1L-1. Nevertheless, the task preserves symmetry between short and long distances, since distances of length $L-\alpha$ and $L+\alpha$ occur with equal frequency. Therefore, the statistical properties of the task do not inherently favor the memorization of earlier study items (i.e., long input-output dependencies).
@@
To build test sequences, the held-out study items were randomly ordered, and queries were generated by first shuffling and then cyclically shifting them (e.g., $(2,8,11,29)\allowbreak\mapsto\allowbreak\{(2,8,11,29),\allowbreak(8,11,29,2),\allowbreak(11,29,2,8),\allowbreak(29,2,8,11)\}$). This design ensured that each study item was queried in all $L$ possible positions. Finally, either the even- or odd-indexed query positions were replaced with random distractors, resulting in a total of $1024\times L\times 2$ test sequences per trial.
In Appendix B, the primacy effect is examined in a more advanced task—associative recall—which has been established as a useful benchmark for evaluating the performance of language model architectures (Olsson et al.,, 2022; Fu et al.,, 2023; Gu and Dao,, 2024).